# From Code to Correctness: Closing the Last Mile of Code Generation with Hierarchical Debugging

### v3

@@ 3.6. Generalization to Real-World Software Defects (RQ5)
While our primary focus is to explore LLMs’ ability to fix LLM-generated code, we also evaluate MGDebugger on repairing real-world, human-written bugs. We leverage the widely used Defects4J dataset (Just et al., 2014), which contains real-world defects from open-source Java projects. We compare MGDebugger with state-of-the-art program repair approaches, including ChatRepair (Xia and Zhang, 2024), Mulpor (Lin et al., 2024), ContrastRepair (Kong et al., 2025), ThinkRepair (Yin et al., 2024), GAMMA (Zhang et al., 2024a), and CodexRepair (Xia et al., 2023). Following prior APR tools (Kong et al., 2025; Xia and Zhang, 2024; Lin et al., 2024; Yin et al., 2024), we use the 255 single-function bugs in Defects4J v1.2 and 82 single-line bugs in Defects4J v2.0 for evaluation and provide the perfect fault localization for each bug. To ensure a fair comparison, we use gpt-3.5-turbo-0301 as the LLM backbone and set the maximum number of generated patches to 200 following prior works (Kong et al., 2025; Xia and Zhang, 2024). For patch validation, we follow the established methodology from prior works (Kong et al., 2025; Xia and Zhang, 2024; Lin et al., 2024; Yin et al., 2024; Zhang et al., 2024a) and conduct manual validation by three experienced engineers who independently assess each patch. A ”correct” patch is defined as one that is semantically equivalent to the ground truth patch, while a ”plausible” patch passes all available test cases but may not fix the underlying bug. The engineers discuss inconsistencies and reach consensus through detailed examination of patch semantics and equivalence to ground truth solutions.
@@
Figure 5 illustrates the overlap of bugs fixed by MGDebugger and four strong baseline approaches (using ChatRepair results from the replication package of ContrastRepair (Kong et al., 2025) due to unavailable bug id details in the original paper (Xia and Zhang, 2024)). The diagram shows 36 bugs fixed by all methods, while MGDebugger uniquely fixes 25 distinct bugs, demonstrating its capability with complex defects. This complementarity with other approaches (e.g., Mulpor’s 26 uniquely fixed bugs) suggests potential for combining methods to achieve even higher repair rates, as different debugging strategies appear effective for various bug types and code contexts.
@@ 4.2. Individual Component Analysis
To further validate the quality of individual components, we conducted manual verification on 38 examples from HumanEval, including 29 successfully debugged and 9 failed cases, using DeepSeek-Coder-V2-Lite. Our analysis reveals three key findings: 1) Code Decomposition: All 38 decomposed versions are functionally equivalent to the original functions, demonstrating 100% semantic correctness in our decomposition process. Since the overall codes have been provided, it’s less likely that the model includes new bugs when transforming the code into subfunctions, demonstrating the high quality of our LLM-based decomposition approach. 2) Test Case Generation: 291 out of 312 generated test cases, representing 93.2% accuracy, were correct when the model is instructed to strictly follow the public test cases when generating test cases for each subfunction. Since our approach derives test cases exclusively from the original public test cases, with a mean test coverage of 71.3%, mutation testing scores of 9.2% and test diversity of 28.6%, we maintain the inherent quality characteristics of the original test suite while enhancing debugging effectiveness. Additionally, in a supplementary experiment on 9 failed debugging cases, providing additional test cases that exposed errors enabled successful debugging of 2 additional functions, indicating potential for further improvement with enhanced test diversity and coverage. 3) Simulated Execution: The LLM correctly predicted 389 of 413 breakpoints, achieving 94.2% accuracy, with failures occurring only in cases requiring complex multi-step reasoning. The high accuracy aligns well with existing studies (Zheng et al., 2024; Ni et al., 2024). Notably, when we manually corrected the prediction errors, 2 of the 9 failed debugging cases could be successfully fixed, indicating potential for further improvement with enhanced simulation accuracy. According to our manual analysis, the high quality of each component in MGDebugger contributes to the overall effectiveness, as demonstrated by the ablation results. This suggests our approach is robust and that further improvements in component quality could lead to even better results.
@@ 4.4. Evaluation on Unseen Data
Since HumanEval and Defects4J are potentially contained in training data of LLMs, potential data leakage could affect our evaluation. To mitigate this threat, we have employed several strategies: (1) we utilized the HumanEvalFix dataset, which consists of human-crafted bugs created after the training cutoff dates of Codestral; (2) our experimental design inherently reduces the impact of potential data leakage by focusing exclusively on test cases where the No-Debugging LLM fails to produce correct solutions (Olausson et al., 2023; Chen et al., 2023b). If data leakage had a strong effect for a particular case, the model would likely have generated a correct solution on its first attempt, and such cases would not be included in our debugging evaluation set.
To further address potential data leakage concerns, we conducted additional experiments on LiveCodeBench (Jain et al., 2025), which contains programming problems released after LLM training cutoffs. Using 131 problems from the February-May 2025 evaluation window with GPT-4o, MGDebugger achieves 36.6% accuracy with 10.8% repair success rate, outperforming all baselines by +7.6% over No-Debugging and +3.0% over Reflexion (Table 6). These results also confirm that MGDebugger’s improvements are not artifacts of data leakage but represent genuine debugging capabilities on completely unseen data.
@@ 4.5. Threats to Validity
While our evaluation is comprehensive, several threats to validity should be acknowledged. 1) The first is the risk of limited generalizability across programming languages and domains. To mitigate this concern, we have evaluated MGDebugger across four distinct benchmarks spanning both Python and Java languages. This comprehensive validation across different language contexts strengthens our confidence in the approach’s broad applicability. 2) Our method’s performance may vary with different LLM architectures. To address this concern, we conducted experiments with various open-source LLMs of different sizes as well as several commercial models. The consistent improvements observed across this diverse set of models suggest that MGDebugger’s benefits are robust across different model architectures and capabilities. 3) For our evaluation on Defects4J, an internal threat arises from the manual validation process used to determine the correctness of the generated patches. To mitigate this concern, we conducted thorough examinations of each patch, following established validation practices from prior work in automated program repair (Xia and Zhang, 2024; Kong et al., 2025; Lin et al., 2024; Yin et al., 2024; Zhang et al., 2024a).
@@ 5.1. Large Language Models for Code
Recent models such as GPT4 (OpenAI, 2023), Codestral (Mistral AI team, 2024), and DeepSeek-Coder (AI, 2024) have advanced code generation through instruction tuning and reinforcement learning from human feedback (RLHF) with mixed code and natural language data (Ziegler et al., 2020; Rafailov et al., 2023; Anthropic, 2024). Code generation with LLMs has been enhanced by various techniques. Some approaches focus on improving the quality of generated code using planning algorithms, transitioning from outlines to detailed implementations (Zhang et al., 2022; Yao et al., 2023; Zelikman et al., 2023; Zheng et al., 2023). Other methods sample multiple programs from the same LLM and rank them to identify the best one (Chen et al., 2023a, 2022; Ni et al., 2023). Additionally, some works leverage multi-agent collaboration frameworks to enhance code generation quality (Zhang et al., 2024b; Huang et al., 2023a; Dong et al., 2024). These approaches aim to optimize the production of correct code from the outset. By contrast, MGDebugger targets the post-generation phase, focusing on debugging and fixing errors that inevitably arise during the code generation process.
@@ 5.2. Repairing LLM-Generated Code
Program repair is a critical aspect of software development, aiming to automatically identify and fix bugs in code (Just et al., 2014; Gupta et al., 2020; Yasunaga and Liang, 2021; Shao et al., 2025). There are two main streams of research in repairing code generated by LLMs: (1) training models to repair code (Huang et al., 2023b; Jiang et al., 2024; Ding et al., 2024; Zheng et al., 2024; Moon et al., 2024; Kumar et al., 2024) and (2) providing external feedback to the raw pretrained models to fix code (Jiang et al., 2023b; Chen et al., 2023b; Olausson et al., 2023; Zhong et al., 2024; Hu et al., 2024). By contrast to previous work that trains separate models for code repair (Ding et al., 2024; Zheng et al., 2024; Moon et al., 2024), MGDebugger does not require task-specific retraining but takes advantage of the inherent capabilities of pretrained LLMs. This flexibility allows MGDebugger to operate in zero-shot settings, offering a lightweight and scalable alternative. Furthermore, exploring the ability of LLMs to fix their own code is a promising direction for self-improvement training of the LLMs (Burns et al., 2023).
MGDebugger falls under the category of work that leverages pretrained models to fix code by reasoning with external feedback. Several recent methods (Zhang et al., 2023; Olausson et al., 2023; Bouzenia et al., 2024; Lee et al., 2024) utilize execution results from test cases to guide LLMs in code correction. More recent works have explored advanced debugging techniques utilizing LLM’s reasoning ability. Reflexion (Shinn et al., 2023) prompts LLMs to reflect on the generated code and uses a memory buffer for iterative refinement. Self-Debugging (Chen et al., 2023b) prompts LLMs to explain or dry run generated programs, known as rubber duck debugging. LDB (Zhong et al., 2024) segments programs into basic blocks, tracking variable values during runtime after each block to verify the correctness against the task description. Although these methods incorporate detailed execution feedback and iterative refinement, they treat the whole function as a single unit and perform sequential debugging, limiting their effectiveness with complex code (Chen et al., 2023b; Hossain et al., 2024). MGDebugger addresses this issue by introducing a hierarchical approach, debugging from low-level errors to high-level flaws. This method ensures a more systematic and accurate debugging process, especially for complex and multifunctional systems.
@@ 5.3. Automated Program Repair
Traditional APR techniques can be broadly classified into three categories: template-based (Liu et al., 2019; Zhang et al., 2024a), heuristic-based (Le Goues et al., 2012; Kim et al., 2024) and constraint based (Le et al., 2017; Mechtaev et al., 2016; Zhu et al., 2023). While these approaches have demonstrated effectiveness, they often face challenges in scalability and patch diversity. The advent of deep learning has shifted the focus to learning-based approaches, particularly those framing program repair as a neural machine translation task (Lin et al., 2024; Li et al., 2020; Lutellier et al., 2020; Jiang et al., 2023a). These methods use encoder-decoder architectures to translate buggy code into correct versions, achieving notable success but remaining limited by the quality and quantity of training data (Yin et al., 2024). More recently, LLM-based APR tools, powered by models like ChatGPT and Codex, have achieved superior fix rates on standard benchmarks like Defects4J (Just et al., 2014) by generating diverse and contextually relevant patches (Xia et al., 2023; Xia and Zhang, 2024; Kong et al., 2025; Yin et al., 2024; Hossain et al., 2024; Xin et al., 2024; Zhang et al., 2025).

### v4

@@ 3.6. Generalization to Real-World Software Defects (RQ5)
While our primary focus is to explore LLMs’ ability to fix LLM-generated code, we also evaluate MGDebugger on repairing real-world, human-written bugs. We leverage the widely used Defects4J dataset (just2014defects4j), which contains real-world defects from open-source Java projects. We compare MGDebugger with state-of-the-art program repair approaches, including ChatRepair (xia2024automated), Mulpor (lin2024one), ContrastRepair (kong2025contrastrepair), ThinkRepair (yin2024thinkrepair), GAMMA (zhang2024gamma), and CodexRepair (xia2023automated). Following prior APR tools (kong2025contrastrepair; xia2024automated; lin2024one; yin2024thinkrepair), we use the 255 single-function bugs in Defects4J v1.2 and 82 single-line bugs in Defects4J v2.0 for evaluation and provide the perfect fault localization for each bug. To ensure a fair comparison, we use gpt-3.5-turbo-0301 as the LLM backbone and set the maximum number of generated patches to 200 following prior works (kong2025contrastrepair; xia2024automated). For patch validation, we follow the established methodology from prior works (kong2025contrastrepair; xia2024automated; lin2024one; yin2024thinkrepair; zhang2024gamma) and conduct manual validation by three experienced engineers who independently assess each patch. A ”correct” patch is defined as one that is semantically equivalent to the ground truth patch, while a ”plausible” patch passes all available test cases but may not fix the underlying bug. The engineers discuss inconsistencies and reach consensus through detailed examination of patch semantics and equivalence to ground truth solutions.
@@
Figure 5 illustrates the overlap of bugs fixed by MGDebugger and four strong baseline approaches (using ChatRepair results from the replication package of ContrastRepair (kong2025contrastrepair) due to unavailable bug id details in the original paper (xia2024automated)). The diagram shows 36 bugs fixed by all methods, while MGDebugger uniquely fixes 25 distinct bugs, demonstrating its capability with complex defects. This complementarity with other approaches (e.g., Mulpor’s 26 uniquely fixed bugs) suggests potential for combining methods to achieve even higher repair rates, as different debugging strategies appear effective for various bug types and code contexts.
@@ 4.2. Individual Component Analysis
To further validate the quality of individual components, we conducted manual verification on 38 examples from HumanEval, including 29 successfully debugged and 9 failed cases, using DeepSeek-Coder-V2-Lite. Our analysis reveals three key findings: 1) Code Decomposition: All 38 decomposed versions are functionally equivalent to the original functions, demonstrating 100% semantic correctness in our decomposition process. Since the overall codes have been provided, it’s less likely that the model includes new bugs when transforming the code into subfunctions, demonstrating the high quality of our LLM-based decomposition approach. 2) Test Case Generation: 291 out of 312 generated test cases, representing 93.2% accuracy, were correct when the model is instructed to strictly follow the public test cases when generating test cases for each subfunction. Since our approach derives test cases exclusively from the original public test cases, with a mean test coverage of 71.3%, mutation testing scores of 9.2% and test diversity of 28.6%, we maintain the inherent quality characteristics of the original test suite while enhancing debugging effectiveness. Additionally, in a supplementary experiment on 9 failed debugging cases, providing additional test cases that exposed errors enabled successful debugging of 2 additional functions, indicating potential for further improvement with enhanced test diversity and coverage. 3) Simulated Execution: The LLM correctly predicted 389 of 413 breakpoints, achieving 94.2% accuracy, with failures occurring only in cases requiring complex multi-step reasoning. The high accuracy aligns well with existing studies (zheng2024opencodeinterpreter; ni2024next). Notably, when we manually corrected the prediction errors, 2 of the 9 failed debugging cases could be successfully fixed, indicating potential for further improvement with enhanced simulation accuracy. According to our manual analysis, the high quality of each component in MGDebugger contributes to the overall effectiveness, as demonstrated by the ablation results. This suggests our approach is robust and that further improvements in component quality could lead to even better results.
@@ 4.4. Evaluation on Unseen Data
Since HumanEval and Defects4J are potentially contained in training data of LLMs, potential data leakage could affect our evaluation. To mitigate this threat, we have employed several strategies: (1) we utilized the HumanEvalFix dataset, which consists of human-crafted bugs created after the training cutoff dates of Codestral; (2) our experimental design inherently reduces the impact of potential data leakage by focusing exclusively on test cases where the No-Debugging LLM fails to produce correct solutions (olausson2023selfrepair; chen2023teaching). If data leakage had a strong effect for a particular case, the model would likely have generated a correct solution on its first attempt, and such cases would not be included in our debugging evaluation set.
To further address potential data leakage concerns, we conducted additional experiments on LiveCodeBench (jain2025livecodebench), which contains programming problems released after LLM training cutoffs. Using 131 problems from the February-May 2025 evaluation window with GPT-4o, MGDebugger achieves 36.6% accuracy with 10.8% repair success rate, outperforming all baselines by +7.6% over No-Debugging and +3.0% over Reflexion (Table 6). These results also confirm that MGDebugger’s improvements are not artifacts of data leakage but represent genuine debugging capabilities on completely unseen data.
@@ 4.5. Threats to Validity
While our evaluation is comprehensive, several threats to validity should be acknowledged. 1) The first is the risk of limited generalizability across programming languages and domains. To mitigate this concern, we have evaluated MGDebugger across four distinct benchmarks spanning both Python and Java languages. This comprehensive validation across different language contexts strengthens our confidence in the approach’s broad applicability. 2) Our method’s performance may vary with different LLM architectures. To address this concern, we conducted experiments with various open-source LLMs of different sizes as well as several commercial models. The consistent improvements observed across this diverse set of models suggest that MGDebugger’s benefits are robust across different model architectures and capabilities. 3) For our evaluation on Defects4J, an internal threat arises from the manual validation process used to determine the correctness of the generated patches. To mitigate this concern, we conducted thorough examinations of each patch, following established validation practices from prior work in automated program repair (xia2024automated; kong2025contrastrepair; lin2024one; yin2024thinkrepair; zhang2024gamma).
@@ 5.1. Large Language Models for Code
Recent models such as GPT4 (openai2023gpt4), Codestral (mistralaiteam2024codestral), and DeepSeek-Coder (zhu2024deepseekcoderv2) have advanced code generation through instruction tuning and reinforcement learning from human feedback (RLHF) with mixed code and natural language data (ziegler2020finetuning; rafailov2023direct; anthropic2024introducing). Code generation with LLMs has been enhanced by various techniques. Some approaches focus on improving the quality of generated code using planning algorithms, transitioning from outlines to detailed implementations (zhang2022planning; zelikman2023parsel; zheng2023outline). Other methods sample multiple programs from the same LLM and rank them to identify the best one (chen2023universal; chen2022codet; ni2023lever). Additionally, some works leverage multi-agent collaboration frameworks to enhance code generation quality (zhang2024codeagent; huang2023agentcoder; dong2024selfcollaboration). These approaches aim to optimize the production of correct code from the outset. By contrast, MGDebugger targets the post-generation phase, focusing on debugging and fixing errors that inevitably arise during the code generation process.
@@ 5.2. Repairing LLM-Generated Code
Program repair is a critical aspect of software development, aiming to automatically identify and fix bugs in code (just2014defects4j; gupta2020synthesize; yasunaga2021breakitfixit; shao2025llms). There are two main streams of research in repairing code generated by LLMs: (1) training models to repair code (huang2023empirical; jiang2024training; ding2024cycle; zheng2024opencodeinterpreter; moon2024coffee; kumar2024training) and (2) providing external feedback to the raw pretrained models to fix code (jiang2023selfevolve; chen2023teaching; olausson2023selfrepair; zhong2024debug; hu2024leveraging). By contrast to previous work that trains separate models for code repair (ding2024cycle; zheng2024opencodeinterpreter; moon2024coffee), MGDebugger does not require task-specific retraining but takes advantage of the inherent capabilities of pretrained LLMs. This flexibility allows MGDebugger to operate in zero-shot settings, offering a lightweight and scalable alternative. Furthermore, exploring the ability of LLMs to fix their own code is a promising direction for self-improvement training of the LLMs (burns2023weaktostrong).
MGDebugger falls under the category of work that leverages pretrained models to fix code by reasoning with external feedback. Several recent methods (zhang2023selfedit; olausson2023selfrepair; bouzenia2024repairagent; lee2024unified) utilize execution results from test cases to guide LLMs in code correction. More recent works have explored advanced debugging techniques utilizing LLM’s reasoning ability. Reflexion (shinn2023reflexion) prompts LLMs to reflect on the generated code and uses a memory buffer for iterative refinement. Self-Debugging (chen2023teaching) prompts LLMs to explain or dry run generated programs, known as rubber duck debugging. LDB (zhong2024debug) segments programs into basic blocks, tracking variable values during runtime after each block to verify the correctness against the task description. Although these methods incorporate detailed execution feedback and iterative refinement, they treat the whole function as a single unit and perform sequential debugging, limiting their effectiveness with complex code (chen2023teaching; hossain2024deep). MGDebugger addresses this issue by introducing a hierarchical approach, debugging from low-level errors to high-level flaws. This method ensures a more systematic and accurate debugging process, especially for complex and multifunctional systems.
@@ 5.3. Automated Program Repair
Traditional APR techniques can be broadly classified into three categories: template-based (liu2019avatar; zhang2024gamma), heuristic-based (legoues2012genprog; kim2024enhancing) and constraint based (le2017s3; mechtaev2016angelix; zhu2023tare). While these approaches have demonstrated effectiveness, they often face challenges in scalability and patch diversity. The advent of deep learning has shifted the focus to learning-based approaches, particularly those framing program repair as a neural machine translation task (lin2024one; li2020dlfix; lutellier2020coconut; jiang2023knod). These methods use encoder-decoder architectures to translate buggy code into correct versions, achieving notable success but remaining limited by the quality and quantity of training data (yin2024thinkrepair). More recently, LLM-based APR tools, powered by models like ChatGPT and Codex, have achieved superior fix rates on standard benchmarks like Defects4J (just2014defects4j) by generating diverse and contextually relevant patches (xia2023automated; xia2024automated; kong2025contrastrepair; yin2024thinkrepair; hossain2024deep; xin2024detecting; zhang2025patch).
@@ 6. Conclusion
Acknowledgments
This research is funded by the National Key Research and Development Program of China (Grant No. 2023YFB4503802) and the Natural Science Foundation of Shanghai (Grant No. 25ZR1401175). The authors would like to thank National Supercomputer Center in Guangzhou for providing high performance computational resources.